# Sphere Shaping with Dispersion-Aware Sequence Selection for Long-Haul Transmission Systems

### 21 Dec 2023 v1

@@ Abstract
We introduce a nonlinear-tolerant design approach to Enumerative Sphere Shaping (ESS) for $M$-QAM signaling, utilizing sequence selection techniques based on a sign-independent metric: the Energy Dispersion Index, EDI (a scheme denoted E-ESS) and a novel sign-dependent metric: EDI of a dispersed sequence, D-EDI (a scheme denoted D-E-ESS). These approaches are designed to minimize rate loss and enhance transmission performance in nonlinear optical fiber transmission systems, catering to both short-distance and long-haul scenarios. Our simulation results reveal significant performance gains over conventional ESS, with improvements up to $0.4$ bits/4D-symbol. These improvements were observed over a 205 km single-span standard single mode fiber link in WDM transmission, with five dual-polarization channels, each operating at a net rate of $400$ Gbit/s. Furthermore, we demonstrate that D-E-ESS surpasses conventional ESS by $0.03$ bits/4D-symbol in achievable information rate over a $30$ span link with $80$ km span in a single-wavelength 8 discrete multi-bands (DMB) transmission with an $880$ Gbit/s net rate per channel, achieving similar performance to sequence selection based on a full split-step Fourier method (SSFM) simulation. We also demonstrate that D-EDI exhibits a more accurate negative correlation with the transmission performance compared to EDI. Notably, our proposed D-E-ESS scheme maintains robust performance improvements even when we consider reduced-complexity versions of it, consistently delivering throughput enhancements across various block lengths and selected sequence lengths.
@@ I Introduction
Numerous studies have shown that achieving high throughput in optical fiber transmission systems is feasible through the utilization of techniques like Polarization Division Multiplexing (PDM), combined with $M$-ary Quadrature Amplitude Modulation (QAM) formats and Forward Error Correction (FEC) codes [1, 2, 3]. However, this increased efficiency often comes at the cost of reduced transmission distances due to effects such as OSNR limitation and Kerr-induced nonlinear distortions [4]. To approach the Shannon limit over an Additive White Gaussian Noise (AWGN) channel, the concept of constellation shaping has emerged as a promising strategy. Kschischang and Pasupathy’s work demonstrated that the probability distribution that maximizes the achievable information rate over an AWGN channel, while maintaining a fixed average energy constraint, corresponds to the Maxwell-Boltzmann (MB) distribution [5, 6, 7].
@@
To effectively implement the MB distribution, various Distribution Matching (DM) techniques have been proposed for Probabilistic Constellation Shaping (PCS). Notable methods include enumerative sphere shaping (ESS) [8] and constant composition distribution matching (CCDM) [9]. CCDM [9] has emerged as a competitive technique for implementing PCS with rectangular or square $M$-QAM constellations by delivering near-optimal linear shaping gains when large amplitude blocks are considered [10]. However, in the context of mitigating Kerr-induced nonlinear interference (NLI) in optical fiber transmission, long block length DM may exhibit a higher susceptibility to performance degradation, as highlighted in previous analyses [11, 12, 13]. These works showed that short block length DM can effectively reduce NLI in optical fiber transmission. Yet, it is crucial to consider that, as the block length decreases, the rate loss increases. Hence, finding a high-performance nonlinear tolerant shaping scheme involves striking a trade-off between nonlinear gain and rate loss [14]. Besides, simulation results in [15, 16] showed that carrier phase recovery (CPR) in the DSP chain at the receiver side can exhibit comparable effectiveness to finite-length PCS in mitigating NLI.
@@
To further enhance the nonlinear performance of PCS schemes, recent research efforts focused on introducing constraints in the trellis grid tree of ESS. For instance, in [17], the concept of kurtosis-limited ESS (K-ESS) was introduced as an extension of ESS. K-ESS not only has a maximum energy limit but also imposes a maximum kurtosis limit on the output sequence. This leads to a reduced average one-dimensional (1D) amplitude kurtosis, which in turn offers an additional nonlinear gain. This effect is underpinned by the EGN model [18], which forecasts a more significant nonlinear penalty associated with higher kurtosis levels. Subsequently, in [19], Band-Limited ESS (B-ESS) was introduced, which uses a specific subset of the trellis tree to limit 1D amplitude variance and kurtosis. B-ESS demonstrated superior performance compared to both K-ESS and conventional ESS. While both K-ESS and B-ESS implement sequences in four-dimensional (4D) space while applying energy limitations over each separate dimension, we proposed in a previous work [20] a band-limited version of ESS with energy limitations in 4D (two pairs of I and Q signals), achieving further throughput enhancement. However, for these three schemes, the rate loss was still significant.
@@
Other design approaches of nonlinear tolerant signaling schemes consist in manipulating the input signal to the DM to generate multiple candidate sequences. Then, specific metrics are used to select “good” sequences, i.e., sequences generating low NLI. For example, the energy dispersion index (EDI), introduced in [21], is a metric that quantifies energy changes of the signal over a given temporal window. EDI showed a negative correlation with the effective Signal-to-Noise Ratio (SNR) of the CCDM-based transmission; however, this negative correlation is observed in a single-polarization transmission scenario. As for ESS-based schemes, our study in [20] highlighted that the EDI, when calculated over 4D-energy levels, demonstrates a notable negative correlation for conventional ESS, B-ESS [19], and 4D-BL-ESS over $110$ Gbaud, $400$ km, $5$ WDM transmission. However, this correlation appears to be inaccurate in the case of K-ESS [17]. In a different work [22], EDI was utilized for sequence selection in a scheme called List-Encoding CCDM (L-CCDM), achieving nonlinear gains with respect to conventional CCDM. Yet, for long-haul transmissions, CCDM is not the best distribution matcher, as similar nonlinear gains can be achieved through conventional ESS with shorter block lengths [14]. More recently, yet another approach, presented in [23], used EDI and a perturbation-model-based low-pass-filtered symbol-amplitude sequence (LSAS) metric [24] for sequence selection. These methods have substantially improved the nonlinear tolerance of ESS or CCDM. However, it is essential to highlight that most significant gains in ESS schemes have been observed over short-distance single-span transmissions.
@@
The concept of sequence selection (SS) was first introduced in [25] and has been further studied in [26]. This innovative signaling approach provided valuable insights into the computation of a lower bound of the optical fiber channel capacity. In particular, the analysis clearly showed that there is considerable room for improvement in optical fiber transmission, especially for long-haul links. A recent study  [27] from the same authors emphasized the importance of enhancing performance through sequence selection using sign-dependent metrics in long-distance transmission scenarios. The authors demonstrated that sign-independent metrics such as EDI, windowed kurtosis, or LSAS do not yield significant throughput improvements over long-haul transmissions. They also showed that a sequence selection based on the computation of NLI through a numerical simulation (Split-Step Fourier Method or SSFM) of a noiseless single-channel propagation of sequences achieved throughput enhancements over long-distance transmissions even when an optimized CPR is applied. However, this selection method is not feasible in practice due to the high complexity of the SSFM.
Figure 1: Multi-block FEC-independent D-E-ESS sequence selection implementations. The parity bits from a current FEC-independent D-E-ESS block are allocated to a subset of the signs in the subsequent block.
@@
In this paper, we introduce a novel sign-dependent metric, the EDI of dispersed sequences, named D-EDI, which accounts for the influence of chromatic dispersion throughout the transmission process. Additionally, we show that D-EDI demonstrates a negative correlation with the effective SNR for high-rate multi-span transmission scenarios. Then, we propose two new sequence-selection ESS (SS-ESS) schemes designed for improving nonlinear tolerance: E-ESS, which uses sign-independent EDI sequence selection, and D-E-ESS [^1] For an ESS system employing sequence selection based on either EDI or D-EDI and comprising $n$ cascaded DMs, where each DM includes $\nu$ flipping bits, we name the scheme $E\text{-}ESS_{n}^{\nu}$ for the EDI-based approach, and $D\text{-}E\text{-}ESS_{n}^{\nu}$ for the D-EDI-based approach. The schemes will be detailed in section III., which employs sequence selection based on sign-dependent D-EDI. We thoroughly investigate and optimize the performance of these schemes in both single-span and multi-span transmission scenarios with optimized CPR. Our D-E-ESS scheme demonstrates superior performance compared to conventional ESS across various block lengths and complexity levels, in both single-channel and WDM transmission scenarios. Remarkably, it performs on par with the ideal SSFM-based sequence selection method, however with reduced complexity. The paper is structured as follows: in Section II, we recall some building blocks from recent literature that we reuse in our proposed scheme; then, in Section III, we define the new metric D-EDI and outline the architecture of the E-ESS and D-E-ESS transmitter schemes. Section IV focuses on the performance of a single-span transmission system and gives insights into the optimization of the proposed methods in this context. In Section V, we study the performance over multi-span single-channel and WDM transmission systems with CPR at the receiver side. Finally, Section VI summarizes key findings and conclusions drawn from our investigations.
II Introduction to mapping strategy and Multi-block FEC-independent sequence selection
II-A Mapping Strategy
Distribution matching can be applied across different dimensions, including 1D, 2D, 4D, or even more. Fig. 3 in [28] provides a visual representation of the mapping strategies employed by ESS schemes up to four dimensions. In the case of 1D mapping, 4D symbols are shaped using four independent amplitude sequences, essentially employing one ESS per dimension. In 2D mapping, each polarization, namely X and Y, utilizes a distinct ESS-shaped amplitude sequence. In contrast, 4D mapping employs a single-shaped sequence for a dual-polarization symbol sequence. For example, assuming the amplitude sequence is $[a_{1},a_{2},\cdots,a_{8}]$, the dual-polarization symbol sequence over two time slots employing 4D mapping will be:
$\begin{bmatrix}S_{Pol.X,T_{1}}&S_{Pol.X,T_{2}}\\ S_{Pol.Y,T_{1}}&S_{Pol.Y,T_{2}}\end{bmatrix}=\begin{bmatrix}\pm a_{1}\pm ja_{2}&\pm a_{5}\pm ja_{6}\\ \pm a_{3}\pm ja_{4}&\pm a_{7}\pm ja_{8}\end{bmatrix}$ (1)
This last approach offers the advantage of reducing the likelihood of having equal amplitudes at the same time slot, resulting in less frequent occurrences of high peak power values and leading to lower levels of non-linear distortions. This advantage has been demonstrated in various studies [11, 17, 28]. Hence, in our work, we use 4D mapping of the amplitudes at the output of the distribution matcher.
II-B Multi-Block FEC-Independent Sequence Selection
The goal of the multi-block FEC-independent SS scheme is to reduce the complexity by moving the FEC from before to after the sequence selection. The implementation of multi-block FEC-independent sequence selection was proposed in [27]. As shown in Fig. 1, the parity bits of each FEC-independent D-E-ESS block (a block that will be presented in Fig. 2 of the next section), are assigned to a portion of the signs in the subsequent block. In the process of selecting a specific sequence, these signs are predetermined by the parity bits from the preceding block. These predetermined signs are required for the calculation of a sign-dependent metric. It is important to note that, once the candidate sequences have been generated, their bits become fixed and unalterable. Consequently, if the FEC encoding is applied after the sequence selection (SS) process, it becomes necessary to shift the parity bits to the subsequent block.
III E-ESS and D-E-ESS design
Figure 2: Block diagram of FEC independent D-E-ESS in the PAS transmitter, EDI of dispersed sequences (D-EDI) is determined by averaging the calculated EDI at multiple locations along an ideal dispersive fiber. The dispersion is applied at $1$ sample per symbol. EDI is a special case of D-EDI where $m_{D}=0$. The sign bits are fixed through a multi-block FEC-independent sequence selection process as shown in Fig. 1.
Figure 3: Structure of cascaded distribution matchers (DMs). The output amplitude sequences of $n$ DMs are concatenated to form a longer sequence.
The fundamental concept behind E-ESS draws inspiration from the list-CCDM [22]. Our approach allows for the use of various distribution matchers. In this study, we specifically focus on ESS. In Fig. 2, we plot the encoding process where input bits are fed into a 1D ESS distribution matcher (DM) with block length $l$ that generates $l$ amplitudes. The input bit stream is composed of two parts: $k-\nu$ information bits and $\nu$ prefix flipping bits. Altering these $\nu$ prefix flipping bits can cause pronounced changes in the sequences’ indices during the ESS encoding process. Such changes are reflected as considerable alterations in the amplitudes across the block. Conveniently, these flipping bits are discarded after the decoding process at the receiver side.

### 05 Mar 2024 v2

# Sequence Selection with Dispersion-Aware Metric for Long-Haul Transmission Systems

@@ Abstract
We introduce a novel sign-dependent metric: the energy dispersion index (EDI) of sequences that endured chromatic dispersion, denoted as D-EDI, which exhibits a more accurate opposite variations with the transmission performance compared to the standard EDI metric. Then, by applying D-EDI and EDI to the sequence selection (SS) process, we present two signaling approaches denoted as D-SS and E-SS respectively. These approaches are designed to minimize rate loss and enhance transmission performance in nonlinear optical fiber transmission systems, catering to both short-distance and long-haul scenarios. With enumerative sphere shaping (ESS) as distribution matcher (DM), our simulation results reveal significant performance gains over ESS without sequence selection (SS), with improvements up to $0.4$ bits/4D-symbol. These improvements were observed over a $205$-km single-span standard single mode fiber link in WDM transmission, with five dual-polarization channels, each operating at a net rate of $400$ Gbit/s. Furthermore, we demonstrate that D-SS surpasses ESS without SS by $0.03$ bits/4D-symbol in achievable information rate over a $30\times 80$ km link in a single-wavelength, with 8 discrete multi-band (DMB) transmission, and an $880$ Gbit/s net rate. Notably, our proposed D-SS scheme achieves similar performance to a sequence selection based on a full split-step Fourier method (SSFM) simulation and it consistently delivers throughput enhancements across various block lengths and selected sequence lengths.
@@ I Introduction
Numerous studies have shown that achieving a high throughput in optical fiber transmission systems is feasible through the utilization of polarization division multiplexing (PDM), combined with $M$-ary quadrature amplitude modulation (QAM) formats and forward error correction (FEC) codes [1, 2, 3]. However, this increased efficiency often comes at the cost of reduced transmission distances due to OSNR limitation and Kerr-induced nonlinear distortions [4]. To approach the Shannon limit over a channel with additive white Gaussian noise (AWGN), constellation shaping has emerged as a promising strategy. Kschischang and Pasupathy’s work demonstrated that the probability distribution that maximizes the achievable information rate over an AWGN channel, while maintaining a fixed average energy constraint, corresponds to the Maxwell-Boltzmann (MB) distribution [5, 6, 7].
@@
To effectively implement the MB distribution, various distribution matching (DM) techniques have been proposed for probabilistic constellation shaping (PCS). Notable methods include enumerative sphere shaping (ESS) [8] and constant composition distribution matching (CCDM) [9]. CCDM [9] has emerged as a competitive technique for implementing PCS with $M$-QAM constellations by delivering near-optimal linear shaping gains when large amplitude blocks are considered [10]. However, in the context of mitigating Kerr-induced nonlinear interference (NLI) in optical fiber transmission, long-block-length DM may exhibit a higher susceptibility to performance degradation, as highlighted in previous analyses. In [11, 12, 13], it was shown that short-block-length DM can effectively reduce NLI. Yet, it is crucial to recall that, as the block length decreases, the rate loss increases. Hence, finding a high-performance nonlinear-tolerant shaping scheme involves striking a trade-off between nonlinear gain and rate loss [14]. In this respect, ESS is more favored compared to CCDM due to its lower rate loss at the same short block lengths. Furthermore, simulation results in [15, 16] showed that carrier phase recovery (CPR) in the DSP chain at the receiver side can exhibit comparable effectiveness to short-length PCS in mitigating NLI.
@@
To further enhance the nonlinear performance of PCS schemes, recent research efforts focused on modifying the DM structure. For ESS, the nonlinear tolerance of its output amplitude sequence was improved by introducing constraints in the grid tree of the ESS, such as the kurtosis-limited ESS (K-ESS) in [17] and the band-Limited ESS (B-ESS) [18]. In both approaches, 1D amplitude variance and/or kurtosis were limited, wich offers a nonlinear gain. This effect is underpinned by the EGN model [19] that forecasts a more significant nonlinear penalty associated with higher kurtosis levels. While both K-ESS and B-ESS implement sequences in four-dimensional (4D) real-valued space while applying energy limitations over each dimension separately, we proposed in [20] a band-limited version of ESS with energy limitations in 4D (two pairs of I and Q signals), achieving higher throughput. However, for these three schemes, the rate loss was still significant.
@@
Other design approaches of nonlinear tolerant signaling schemes consist of manipulating the input signal to the DM to generate multiple candidate sequences. Then, specific metrics are used to select “good” sequences, i.e. sequences that generate low NLI levels. For example, the energy dispersion index (EDI), introduced in [21], is a metric that quantifies energy changes of the signal over a given temporal window. EDI showed an opposite variation with the effective signal-to-noise ratio (SNR) of the CCDM-based transmission. However, this behavior was observed in a single-polarization transmission scenario. As for ESS-based schemes, our study in [20] highlighted that the EDI, when calculated over 4D-energy levels, demonstrates a notable opposite variation with SNR for ESS, B-ESS [18], and 4D-BL-ESS over $110$ Gbaud, $400$ km, $5$ WDM transmission. However, this behavior appeared to be inaccurate in the case of K-ESS [17]. In a different work [22], EDI was utilized for sequence selection in a scheme called list-encoding CCDM (L-CCDM), achieving nonlinear gains with respect to conventional CCDM. Yet, for long-haul transmissions, CCDM is not the best distribution matcher, as similar nonlinear gains can be achieved through ESS with shorter block lengths [14]. More recently, yet another approach [23] used EDI and a perturbation-model-based low-pass-filtered symbol-amplitude sequence (LSAS) metric [24] for sequence selection. These methods have substantially improved the nonlinear tolerance of ESS or CCDM. However, it is essential to highlight that the most significant gains in ESS schemes have been observed over short-distance single-span transmissions.
@@
The concept of sequence selection (SS) was first introduced in [25] and has been further studied in [26]. This innovative signaling approach provided valuable insights into the computation of a lower bound of the optical fiber channel capacity. In particular, the analysis clearly showed that there is considerable room for improvement in optical fiber transmission, especially for long-haul links. A recent study  [27] from the same authors emphasized the importance of performing sequence selection using sign-dependent metrics in long-haul transmission scenarios. The authors demonstrated that sign-independent metrics such as EDI, windowed kurtosis, or LSAS do not yield significant throughput improvements over long-haul links. They also showed that a sequence selection based on the computation of NLI through a numerical simulation (split-step Fourier method or SSFM) of a noiseless single-channel propagation of sequences achieved throughput enhancements over long-haul transmissions even when an optimized CPR is applied. However, this method is not feasible in practice due to the high complexity of the SSFM.
@@
In this paper, we introduce a novel sign-dependent metric, the EDI of dispersed sequences, named D-EDI, which accounts for the influence of chromatic dispersion throughout transmission. Additionally, we show that D-EDI demonstrates a negative correlation with the effective SNR for high-rate multi-span scenarios. By negative correlation, we mean that the two metrics vary in opposite directions. Next, we utilize EDI and D-EDI as metrics for sequence selection, with ESS serving as the DM. The approach that employs sign-independent EDI for sequence selection is termed E-SS, and the one using sign-dependent D-EDI is designated as D-SS [^1] For a PCS system employing sequence selection based on either EDI or D-EDI and comprising $n$ cascaded DMs, where each DM includes $\nu$ flipping bits, we name the scheme $\mathrm{E\text{-}SS_{n}^{\nu}}$ for the EDI-based approach, and $\mathrm{D\text{-}SS_{n}^{\nu}}$ for the D-EDI-based approach. The schemes will be detailed in section IV.. We thoroughly investigate and optimize the performance of these schemes in both single-span and multi-span transmission scenarios with optimized CPR. Our D-SS scheme demonstrates superior performance compared to ESS without sequence selection across various block lengths and complexity levels, in both single-channel and WDM transmission scenarios. Remarkably, it performs on par with the ideal SSFM-based sequence selection method, however with a reduced complexity. The paper is structured as follows: in Section II, we describe the simulated transmission systems, then, in Section III, we define the new metric D-EDI. In Section IV, we outline the architecture of the E-SS and D-SS transmitter schemes. Section V focuses on the performance of a single-span transmission system and gives insights into the optimization of the proposed methods in this context. In Section VI, we study the performance over multi-span single-channel and WDM transmission systems with CPR at the receiver side. Finally, Section VII summarizes key findings and conclusions drawn from our investigations.
II System Description
We study the performance of E-SS and D-SS in two scenarios: first, a single-span transmission scenario similar to the one in [17, 18]; second, a multi-span long-haul transmission using digital-multi-band (DMB) format. Through the two studies, we observe the effects of nonlinear distortions with various signaling schemes. NLI distortions are dominant at high power levels, therefore, they primarily occur over the first kilometers of the span in a single-span transmission, and occur at the start of each span in a multi-span scenario with discrete optical amplification.
Our simulated single-span scenario consists of a transmission over $205~{}\mathrm{km}$ of standard single-mode fiber (SSMF), to perform a consistent comparison with [17, 18] in which the kurtosis-limited (K-ESS) and band-limited (B-ESS) sphere shaping were introduced. We perform probabilistic shaping using PDM 64-QAM modulation with a block length $l=108$ to align with the conditions described in [17, 18]. We also consider the FEC-independent sequence selection proposed in [27] and ESS [8] for distribution matching. We also use a 4D mapping strategy [28] of our shaped amplitudes which has usually shown better performance than 1D and 2D mapping [23]. Our transmitted signal comprised $5$ wavelength-division multiplexing (WDM) channels, each operating at $50$ GBaud, resulting in a raw data rate of $600$ Gbit/s per channel. For all the tested ESS schemes, we employed a low-density parity-check (LDPC) code with a length of $64800$ bits, following the DVB-S2 standard, and a code rate $r_{c}=5/6$. Consequently, the achieved net bit rate amounted to $8$ bits per 4D symbol, resulting in a net bit rate of $400$ Gbit/s per channel. The channel spacing is set to $55$ GHz, and we apply root-raised cosine (RRC) pulse shaping with a roll-off factor of $0.1$ to each channel. SSMF is simulated with an attenuation coefficient $\alpha_{\mathrm{dB}}=0.2~{}\mathrm{dB/km}$, chromatic dispersion coefficient $D=17~{}\mathrm{ps/nm/km}$, polarization mode dispersion (PMD) of $0.04~{}\mathrm{ps/\sqrt{km}}$, and a nonlinear parameter $\gamma=1.3~{}\mathrm{(W\cdot km)^{-1}}$ at $\lambda=1550~{}\mathrm{nm}$. We do not add any laser phase noise in the simulation to simplify our analysis. Following propagation, the central channel underwent several signal processing steps, including optical filtering, matched filtering, chromatic dispersion compensation, and genie-aided multiple-input-multiple-output (MIMO) channel equalization. Subsequently, for carrier phase recovery (CPR), we applied a fully-data-aided phase filter with a window averaging over $64$ symbols to compensate for phase rotation induced by cross-phase modulation (XPM), as elaborated in [29]. Finally, we measured the electric SNR denoted $\mathrm{SNR_{elec}}$ from the equalized constellations and computed the generalized mutual information (GMI) using Eq. (10) from [30].
For the higher baud rate long-haul link scenario, we evaluate the potential advantages of our E-SS and D-SS schemes employing digital-multi-band (DMB) signals known for their superior resistance to Kerr effects compared to single-carrier schemes [31, 32]. The baud rate in this scenario is set at $110$ Gbaud, with $8$  digital subcarriers, first in a single-wavelength configuration, then in a 5 WDM-channel configuration. The link is made of $30\times 80$ km SSMF spans with one $5$ dB-noise-figure EDFA amplifier per span. The achieved net data rate per channel is $880$ Gbit/s. The digital signal processing (DSP) blocks employed in this scenario are consistent with those detailed in the preceding paragraph.
III From EDI to D-EDI
EDI is a sign-independent metric proposed in [21]. We denote it as $\Psi\left[\mathbf{X}\right]$ in Eq. (1) defined as the ratio of the variance to the mean of the windowed energies in the sequence $\mathbf{G}^{w}$ defined in Eq. (2). For a given window length $w$ and a $2\times L_{s}$ sequence $\mathbf{X}$ defined in Eq. (3) as $L_{s}$ 2D complex-valued column vectors or equivalently two $L_{s}$-long row vectors, each element $G_{i}^{w}$ of $\mathbf{G}^{w}$ is computed as the sum of the energies of $w+1$ 2D symbols centered on $\mathbf{x}_{i}$ as shown in Eq. (4). $\mathrm{Var[\cdot]}$ and $\mathrm{E[\cdot]}$ are the variance and expectation of $\cdot$ respectively.
$\Psi\left[\mathbf{X}\right]\overset{\Delta}{=}\frac{\mathrm{Var}\left[\mathbf{G}^{w}\right]}{\mathrm{E}\left[\mathbf{G}^{w}\right]}$ (1)
$\mathbf{G}^{w}=\left[G_{\frac{w}{2}+1}^{w},G_{\frac{w}{2}+2}^{w},\cdots,G^{w}_{i},\cdots,G_{L_{s}-\frac{w}{2}}^{w}\right]$ (2)
$\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{L_{s}}\right]=\begin{bmatrix}\mathbf{x}_{\mathrm{pol_{1}}}\\ \mathbf{x}_{\mathrm{pol_{2}}}\end{bmatrix}$ (3)
$G_{i}^{w}=\sum_{k=i-\frac{w}{2}}^{i+\frac{w}{2}}\|\mathbf{x}_{k}\|^{2},~{}~{}i:\frac{w}{2}+1\rightarrow L_{s}-\frac{w}{2}$ (4)